OpenStudy (anonymous):

Solve: sinxcosx=root3/4 Please? :)

jimthompson5910 (jim_thompson5910):

hint: use the identity sin(2x) = 2*sin(x)*cos(x)

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\sin(2x)=\sqrt[3]{4}\]

OpenStudy (anonymous):

try to work it out from there

OpenStudy (anonymous):

is it cube root 4 or root of 3/4

OpenStudy (anonymous):

square root of 3 over 4

OpenStudy (tkhunny):

?? \(\sin(x)\cos(x) = \dfrac{\sqrt{3}}{4}\) \(2\sin(x)\cos(x) = \dfrac{\sqrt{3}}{2}\) \(\sin(2x) = \dfrac{\sqrt{3}}{2}\) \(2x = \pi/3\;or\;2\pi/3\) \(x = \pi/6\;or\;\pi/3\)

OpenStudy (anonymous):

Thank you!!

OpenStudy (jmark):

Given sinxcosx=root3/4 divide the equation by cosx \[\frac{sinx cosx}{cosx} = \frac{\sqrt{\frac{3}{4}}}{cosx} {sinx} = \frac{\sqrt{\frac{3}{4}}}{cosx} {sinx} = \frac{\sqrt{3}}{2 cosx} {sinx} - \frac{\sqrt{3}}{2 cosx} = 0\] multiply with denominator \[sinx.2cosx - \frac{ \sqrt{3} }{2 cosx } =0\] \[sinx. 2 cosx -\sqrt{3} = 0\] by making use of sin2x = sinx.2cosx \[\sin2x - \sqrt{3} = 0\] \[\sin2x = \sqrt{3}\]